What is cycloid motion.
1
Expert's answer
2019-12-16T11:08:35-0500

A cycloid is the curve traced by a point on the rim of a circular wheel as the wheel rolls along a straight line without slipping.

The parametric equations that define the motion of a cycloid are


"x=r(t-\sin{t}), y=r(1-\cos{t})"